OpenStudy (anonymous):

HELP PLEASE HELP MEEEEEEEEEEEEEEEE MEDALS+FAN!!!!!!!!!!!!!!!!!!!!!!! Solve using the guess-and-test strategy. The math club has 57 members who are in sixth or seventh grades. There are twice as many sixth graders as seventh graders. How many sixth graders are in the club? A. 38 B. 37 C. 19 D. 18

OpenStudy (anonymous):

38

OpenStudy (jjwalton04):

its a because 38/2=19 and 19+38 is 57

OpenStudy (jjwalton04):

BEWM

OpenStudy (anonymous):

i believe b

OpenStudy (anonymous):

ya b

OpenStudy (jjwalton04):

explain then runninthruthe6

OpenStudy (anonymous):

oh nvm

OpenStudy (anonymous):

thank u I wish I could all u guys a medal

OpenStudy (jjwalton04):

well we can give each other medals

OpenStudy (anonymous):

2y +y = 57 Then solve for y: 3y = 57 y = 19 But remember "y" is the number of 7th graders. So now that we know what "y" is, we can plug that value into the original equation and solve for x (the # of 6th graders). x + 19 = 57 x = 38

OpenStudy (anonymous):

X=2y